# A Unified Framework for Inference with General Missingness Patterns and Machine Learning Imputation

### v1

@@ 4 Numerical Experiments
In this section, we conduct data simulation based on synthetic data to show the operating characteristics of the proposed method to corroborate the theoretical prediction that the proposed approach produces consistent estimation for parameters of interest and yields valid uncertainty quantification. We compare our proposed approach with the following baseline methods: 1) unweighted complete-case analysis, where parameter estimation is performed using only fully observed data (“CCA”); 2) complete-case analysis of inverse probability weighted (IPW), where estimation is carried out using fully observed data weighted by true or estimated propensity scores (“WCCA”); 3) multiple imputation (MI) (Rubin,, 1976), a classical and popular method to deal with dataset with non-monotone missingness pattern under the MAR assumption. In this work, we use the implementation of MI provided by the MICE 3.16.0333https://github.com/amices/mice package in R; 4) PPI++ (Angelopoulos et al., 2023b, ), a recent extension of the prediction-powered inference framework that guarantees better estimation efficiency compared to the CCA approach. For PPI++, both fully observed cases and cases with missing outcomes are used for parameter estimation; cases with missing covariates will be dropped.
@@ 4.1 Experimental setup
where the covariates are generated from $X_{i1}=0.1\exp(Z_{i1})+\tau_{i}$, $X_{i2}=\sin(Z_{i2})+\nu_{i}$, with $(Z_{i1},Z_{i2})$ randomly and independently drawn from a bivariate normal distribution with mean zero and a homogeneous marginal variance $\sigma_{z}^{2}$ and Pearson correlation $\rho$. The error terms are independently distributed as $\epsilon_{i}\sim{\mathcal{N}}(0,\sigma^{2})$, $\tau_{i}\sim{\mathcal{N}}(0,\sigma_{\tau}^{2})$, $\nu_{i}\sim{\tt Exponential}(\lambda)$, respectively. To connect with the notation $\bm{Y}_{i}$ introduced in Section 3, we have $\bm{Y}_{i}=(Y_{i},X_{i1},X_{i2},Z_{i1},Z_{i2})$ in this simulation setting. Both $Z_{i1}$ and $Z_{i2}$ are assumed to be fully observed but are assumed not intended to be adjusted in the analysis model. They are part of the simulation to mimic the scenario in which complete and low-cost external information can be used to produce ML predictions of possibly missing $(Y_{i},X_{i1},X_{i2})$ that are required in an analysis model of scientific interest. We set ${\mathcal{C}}_{i}=\infty$ to indicate that all $Y_{i},X_{i1},X_{i2}$ are observed. We introduce three distinct missingness patterns in addition to the fully observed data as follows: ${\mathcal{C}}_{i}=1$ indicates that $Y_{i}$ is missing; ${\mathcal{C}}_{i}=2$ indicates that only $X_{i2}$ is missing; ${\mathcal{C}}_{i}=3$ indicates that both $Y_{i}$ and $X_{i1}$ are missing. We assume a MAR mechanism. We define a propensity score model in the form proposed by Sun and Tchetgen Tchetgen, (2018) as follows:
@@
The average number of subjects under each missingness pattern in 500 simulations is $1,469.9$ for complete cases, and $1,350.7$, $829.5$, $1,349.9$ for ${\mathcal{C}}=1,2,3$, respectively.
@@
$\widehat{X}_{i1}=X_{i1}+b_{i1}+\epsilon_{i1},$
$\widehat{X}_{i2}=X_{i2}+b_{i2}+\epsilon_{i2},$
$\widehat{Y}_{i}=Y_{i}+b_{i3}+\epsilon_{i3}.$
Here both $\epsilon_{i1}$, $\epsilon_{i2}$, and $\epsilon_{i3}$ follow Normal distribution ${\mathcal{N}}(0,\sigma^{2}_{\text{pred}})$, while $b_{i1}$, $b_{i2}$, and $b_{i3}$ follow Exponential distribution ${\tt Exponential}(\lambda_{\text{pred}})$. The two types of error terms represent the noise and bias imposed on the ML predictions, respectively. When these noises and bias are fixed constants 0, we have the perfect ML predictions.
In our simulation, we fix $\sigma=0.5$, $\sigma_{\tau}=0.3$, $\lambda=0.02$, $\sigma_{z}=0.2$, and $\rho=0.4$. To investigate the impact of the quality of the ML predictions on the performance of our proposed methods, we iterate both $\sigma^{2}_{\text{pred}}$ and $\lambda_{\text{pred}}$ from $\{0,0.2,0.4,0.6,0.8,1.0,1.5,2.0\}$. For propensity scores, we adopt three settings. In the first setting, we assume that the propensity score model is unknown yet correctly specified in terms of both the functional forms and covariates. Under this assumption, we estimate the propensity score by maximizing the likelihood function proposed in Sun and Tchetgen Tchetgen, (2018). In the second setting, we again assume an unknown propensity score model in the form proposed by Sun and Tchetgen Tchetgen, (2018), while we only include $Z_{1}$ in the propensity score model to make it misspecified. This setting is used to investigate the robustness of the PS-PPI procedure against propensity score model misspecification. In the final setting, we assume that the true propensity score is known and can be directly applied to estimate the parameters of interest $\bm{\beta}$}. We set the sample size $N=5,000$ and repeat the simulation for $500$ times.
@@ 4.2 Evaluation metrics
To evaluate the performance of included methods in synthetic data simulation, we compute the point estimates and report three key metrics: 1) the average estimated bias; 2) the coverage rates of the 95% confidence intervals (CIs); 3) the average width of these intervals. Specifically, suppose that $M$ replicates are generated in data simulation, and for the $j$-th replicate, let $\widehat{\beta}^{(j)}$ be the point estimate for a scalar parameter, and $(L^{(j)},U^{(j)})$ its 95% CI. We compute the average estimated bias as follows:
$\texttt{Avg. Estimated Bias}=\frac{1}{M}\sum^{M}_{j=1}\left(\widehat{\beta}^{(j)}-\beta_{\text{true}}\right).$
We compute the coverage rate as follows:
$\texttt{Coverage Rate}=\frac{1}{M}\sum^{M}_{j=1}I\left(\beta_{\text{true}}\in(L^{(j)},U^{(j)})\right).$
Finally, we compute the average width of the 95% CIs as follows:
$\texttt{Avg. 95\% CI Width}=\frac{1}{M}\sum^{M}_{j=1}\left(U^{(j)}-L^{(j)}\right).$
@@ 4.3 Results
Figure 2: Widths of 95% confidence intervals, coverage rates, and biases for different methods under the assumption of unknown propensity scores while with estimated propensity scores from correctly specified propensity score models. The X-axis in all panels represents the prediction noise level used in the simulations. The bias level is fixed at $\lambda_{\text{pred}}=0$. For Y-axis: in (a), shows the $\log_{10}$ of the widths of the 95% confidence intervals; in (b), indicates the coverage rate of the 95% confidence intervals; in (c), shows the bias of the estimated parameters. In (b), the red dashed line represents the predefined confidence level 0.95.
We present the results with varying noise levels and fixed bias level ($\lambda_{\text{pred}}=0$), and estimated propensity scores in Figure 2. Additional experimental results on varying prediction bias, as well as different settings of the propensity score models are provided in Figure S1S2S3S4, and S5, in Section S2.1 of the Supplementary Materials. Note that the coverage rate, the width of the 95% CIs, and the bias remain unchanged for CCA, WCCA, and MI when varying the quality of the ML predictions, as they do not apply these predictions in their estimating procedures.
@@
In terms of the coverage rate of the 95% CIs, the CCA approach achieves a coverage rate of 51.0% and 81.6% for $\beta_{0}$ and $\beta_{1}$, respectively, which is below the nominal coverage rate of 95%. This is expected because it does not incorporate propensity scores to adjust for the missing-at-random mechanism that renders the complete-cases not representative of the entire population. WCCA achieves conservative coverage rates of 98.2%, 98.0%, and 96.2% on the three coefficients, respectively. The conservativeness is due to the use of the robust sandwich variance that does not account for the uncertainty from the estimated propensity score model. From Figure S2, where all the settings are the same except that the propensity score model is known, WCCA exhibits valid coverages around 95% for all the three coefficients. MI achieves permissive coverage rates of 93.2%, 91.6% and 56.6%, respectively for the three parameters, falling short of the nominal 95% level. Under perfect predictions, PPI++ achieves coverage rates of 71.6% and 89.2% for $\beta_{0}$ and $\beta_{1}$ due to the fact that it does not use the propensity score, similar to the CCA approach. When the quality of the ML predictions degrades, the coverage rates of these two coefficients decrease and converge to levels similar to the CCA approach. PS-PPI, on the other hand, achieves a coverage rate of around 95% when the perfect predictions are available. When the quality of the ML predictions decreases, it tends to provide a conservative confidence intervals, similar to its base estimator, i.e., the WCCA estimator.
@@
The widths of the 95% CIs for all the three coefficients obtained from CCA are 0.027, 0.085, and 0.129, respectively; the widths of the 95% CIs for all the three coefficients obtained from WCCA are 0.028, 0.087, and 0.130, respectively. For both CCA and WCCA, the width of the 95% CIs are similar for all the three coefficients. MI has 95% CIs of width 0.025, 0.077, and 0.130 on all the three coefficients (however the real coverage rates for $\beta_{1}$ and $\beta_{2}$ are unsatisfactory). PS-PPI yields uniformly narrower 95% CIs compared to WCCA, the only method with valid coverage rates for all the three coefficients. Furthermore, the widths of 95% CIs obtained from PS-PPI is strongly impacted by the quality of the ML predictions. Specifically, PS-PPI tends to produce narrower 95% CIs when ML predictions have better quality. As a comparison, the PPI method shows similar patterns in terms of the relationship of ML predictions quality and the width of 95% CIs. Furthermore, PS-PPI produces narrower 95% CIs than PPI in most of the cases, except the 95% CI of $\beta_{2}$ when the quality of ML predictions becomes mediocre.
@@
From the bias evaluation, it can be seen that CCA, PPI++, and MI’s failure in achieving 95% nominal coverage are primarily due to its biased estimation. For the bias in the estimation of $\beta_{0}$ and $\beta_{1}$, WCCA results in a simulated average bias of 0.0005 and 0.0006, respectively. On the other hand, PS-PPI results in 0.0007 and 0.0008 with the worst quality of ML predictions. Nevertheless, for the parameter where CCA, PPI++ and MI fails to achieve 95% nominal coverage, they exhibit greater simulated bias. Specifically, MI exhibits a simulated average bias of 0.115 for the estimation of $\beta_{2}$, while CCA results in around 0.043 of simulated bias for $\beta_{1}$. In terms of PPI, the bias increases from 0.023 to 0.043 when the quality of ML predictions degrades.
Similar results are also observed from Figure S1S2S3. The simulation results demonstrate that under non-monotone missingness patterns with MAR assumption for the missing mechanism, the proposed PS-PPI approach can produce consistent estimation and valid inference with efficiency no worse than WCCA when the propensity score model is correctly estimated or known. Such consistency and validity in inference are irrelevant to the quality of the ML predictions. Furthermore, when precise ML predictions are available, PS-PPI achieves greater estimation efficiency compared to WCCA. This corroborates our theoretical results.
@@
The simulation results under misspecified propensity score models are shown in Figure S4 and S5. Note that WCCA and PS-PPI are the only two methods that are impacted by the propensity score model specification. When propensity score model is misspecified in terms of the covariates, all the methods show a degree of performance in the width of the 95% CIs similar to the one shown in Figure 2. In terms of the coverage rate, Figure S4 and S5 show that WCCA fails to achieve the nominal 95% coverage rate for $\beta_{0}$ and $\beta_{1}$. PS-PPI achieve the 95% coverage rate when the ML predictions are sufficiently precise, while these coverage rates decline when the quality of ML predictions degrades. However, it can be seen that the coverage rates from PS-PPI still consistently outperform the WCCA baseline. Furthermore, PS-PPI also maintains a higher coverage rate compared to the PPI method, which only utilizes the fully-observed and missing-outcome subsets of data.

### v2

@@ 4 Numerical Experiments
In this section, we conduct simulation studies to show the operating characteristics of the proposed method to corroborate the theoretical predictions of estimation consistency, inferential validity, and efficiency gains. We compare our proposed approach with the following alternatives: 1) unweighted complete-case analysis (“CCA”); 2) inverse probability weighted (IPW) complete-case analysis (“WCCA”), with the weights set at true or estimated propensity scores; 3) multiple imputation (MI) (rubin1976inference), a classical and popular method to deal with datasets with a non-monotone missingness under the MAR assumption. In this work, we use the implementation of MI provided by the MICE 3.16.0333https://github.com/amices/mice package in R (van2011mice); 4) PPI++ (angelopoulos2023PPI++), a recent extension of the prediction-powered inference framework that guarantees better estimation efficiency compared to the CCA approach. For PPI++, fully observed cases and the cases with missing outcomes are used for parameter estimation; the cases with other missingness patterns will be dropped.
@@ 4.1 Experimental setup
where the covariates are generated from $X_{i1}=0.1\exp(Z_{i1})+\tau_{i}$, $X_{i2}=\sin(Z_{i2})+\nu_{i}$, with $(Z_{i1},Z_{i2})$ randomly and independently drawn from a bivariate normal distribution with mean zero and a homogeneous marginal variance $\sigma_{z}^{2}$ and Pearson correlation $\rho$. The error terms are independently distributed as $\epsilon_{i}\sim{\mathcal{N}}(0,\sigma^{2})$, $\tau_{i}\sim{\mathcal{N}}(0,\sigma_{\tau}^{2})$, $\nu_{i}\sim{\tt Exponential}(\lambda)$, respectively. To connect with the notation $\bm{Y}_{i}$ introduced in Section 3, we have $\bm{Y}_{i}=(Y_{i},X_{i1},X_{i2},Z_{i1},Z_{i2})$ in this simulation setting. $Z_{i1}$ and $Z_{i2}$ are assumed to be fully observed and not used in the main analysis model. They can be used as inputs to produce ML predictions of possibly missing $(Y_{i},X_{i1},X_{i2})$ in the analysis model of scientific interest. We set ${\mathcal{C}}_{i}=\infty$ to indicate that all $Y_{i},X_{i1},X_{i2}$ are observed. We introduce three distinct missingness patterns in addition to the fully observed data as follows: ${\mathcal{C}}_{i}=1$ indicates that $Y_{i}$ is missing; ${\mathcal{C}}_{i}=2$ indicates that only $X_{i2}$ is missing; ${\mathcal{C}}_{i}=3$ indicates that both $Y_{i}$ and $X_{i1}$ are missing. We assume an MAR mechanism. We define a propensity score model in the form proposed by sun2018inverse as follows:
@@
$\widehat{X}_{i1}=X_{i1}+b_{i1}+\epsilon_{i1},\widehat{X}_{i2}=X_{i2}+b_{i2}+\epsilon_{i2},\widehat{Y}_{i}=Y_{i}+b_{i3}+\epsilon_{i3}.$
Here both $\epsilon_{i1}$, $\epsilon_{i2}$, and $\epsilon_{i3}$ follow Normal distribution ${\mathcal{N}}(0,\sigma^{2}_{\text{pred}})$, while $b_{i1}$, $b_{i2}$, and $b_{i3}$ follow Exponential distribution ${\tt Exponential}(\lambda_{\text{pred}})$. The two types of error terms represent the noise and bias imposed on the ML predictions, respectively. When the noise and bias are set to 0, perfect ML predictions result.
In our simulation, we fix $\sigma=0.5$, $\sigma_{\tau}=0.3$, $\lambda=0.02$, $\sigma_{z}=0.2$, and $\rho=0.4$. To investigate the impact of the quality of the ML predictions on the performance of our proposed methods, we iterate both $\sigma^{2}_{\text{pred}}$ and $\lambda_{\text{pred}}$ from $\{0,0.2,0.4,0.6,0.8,1.0,1.5,2.0\}$. For propensity scores, we adopt three settings. In the first setting, we assume that the propensity score model is unknown yet correctly specified in terms of both the functional forms and covariates. Under this assumption, we estimate the propensity score by maximizing the likelihood function proposed in sun2018inverse. In the second setting, we again assume an unknown propensity score model in the form proposed by sun2018inverse, while we only include $Z_{1}$ in the propensity score model to make it misspecified. This setting is used to investigate the robustness of the PS-PPI procedure against propensity score model misspecification. In the final setting, we assume that the true propensity score is known and can be directly applied to estimate the parameters of interest $\bm{\beta}$}. We set the sample size $N=5,000$ and repeat the simulation $500$ times. The average number of subjects under each missingness pattern in the 500 simulations is $1,469.9$ for complete cases, and $1,350.7$, $829.5$, $1,349.9$ for ${\mathcal{C}}=1,2,3$, respectively.
@@ 4.2 Evaluation metrics
To evaluate the performance of the proposed and the four alternative methods in simulations, we report three key metrics: 1) estimation bias; 2) the coverage rates of the 95% confidence intervals (CIs); 3) the average width of these intervals. Specifically, suppose that $M$ replicates are generated in data simulation, and for the $j$-th replicate, let $\widehat{\beta}^{(j)}$ be the point estimate for a scalar parameter, and $(L^{(j)},U^{(j)})$ its 95% CI. We compute the estimation bias by $\texttt{Bias}=\frac{1}{M}\sum^{M}_{j=1}\left(\widehat{\beta}^{(j)}-\beta_{\text{true}}\right).$ We compute the coverage rate by $\texttt{Coverage Rate}=\frac{1}{M}\sum^{M}_{j=1}I\left(\beta_{\text{true}}\in(L^{(j)},U^{(j)})\right).$ Finally, we compute the average width of the 95% CIs by $\texttt{Avg. 95\% CI Width}=\frac{1}{M}\sum^{M}_{j=1}\left(U^{(j)}-L^{(j)}\right).$
@@ 4.3 Results
Figure 2: Performance comparison among the proposed (PS-PPI) and four alternative methods (CCA, WCCA, MI, PPI++). (a): Widths of 95% confidence intervals ($\log_{10}$ transformed); (b): coverage rates of the 95% confidence intervals (red dashed lines indicate 95%); (c): estimation biases for different methods with estimated propensity scores based on correctly specified propensity score models. The X-axis in all the panels represents the prediction noise level used in the simulations. The prediction bias level is fixed at $\lambda_{\text{pred}}=0$.
In the main text, we present the results under varying noise levels with a fixed prediction bias ($\lambda_{\text{pred}}=0$), and estimated propensity scores in Figure 2. Additional experimental results obtained under varying prediction biases, as well as different settings of the propensity score models are provided in Figures S1, S2, S3, S4, and S5, Section S2.1 of the Supplementary Materials. Note that because CCA, WCCA, and MI do not use ML predictions, the coverage rate, the width of the 95% CIs, and the bias remain constant with respect to the quality of the ML predictions.
@@
Figure 2(b) shows that the CCA approach achieves coverage rates of 51.0% and 81.6% for $\beta_{0}$ and $\beta_{1}$, respectively, which are below the nominal level of 95%. This is expected because CCA does not use propensity scores to adjust for the missing-at-random mechanism that renders the complete cases non-representative of the entire population. WCCA achieves conservative coverage rates of 98.2%, 98.0%, and 96.2% on the three coefficients, respectively. The conservativeness is due to the use of the robust sandwich variance that does not account for the uncertainty from the estimated propensity score model. From Figure S2 in the Supplementary Materials, when the propensity score model is known, WCCA exhibits valid coverages around 95% for the three coefficients. MI achieves coverage rates of 93.2%, 91.6% and 56.6%, respectively for $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$, falling short of the nominal 95% level. Under perfect predictions, PPI++ achieves coverage rates of 71.6% and 89.2% for $\beta_{0}$ and $\beta_{1}$ due to the fact that it does not use the propensity score, similar to the CCA approach. When the quality of the ML predictions degrades, the coverage rates for $\beta_{0}$ and $\beta_{1}$ decrease and converge to levels similar to the CCA approach. PS-PPI, on the other hand, achieves a coverage rate of around 95% when the perfect predictions are available and tends to provide conservative confidence intervals when the quality of the ML predictions deteriorates, similar to the behavior of its base WCCA estimator.
@@
Based on Figure 2(a), the widths of the 95% CIs for all three coefficients obtained from CCA are 0.027, 0.085, and 0.129, respectively; the widths for WCCA are similar (0.028, 0.087, and 0.130, respectively). MI has 95% CIs of width 0.025, 0.077, and 0.130 for the three coefficients; however, the empirical coverage rates for $\beta_{1}$ and $\beta_{2}$ are below nominal levels. PS-PPI yields uniformly narrower 95% CIs compared to WCCA, the only alternative method with valid coverage rates for all three coefficients. Furthermore, the widths of 95% CIs obtained from PS-PPI are strongly impacted by the quality of the ML predictions. Specifically, PS-PPI tends to produce narrower 95% CIs when ML predictions have better quality. The PPI++ method shows a similar trend, but has wider 95% CIs than PS-PPI, except for the 95% CI of $\beta_{2}$ when the quality of ML predictions becomes mediocre.
@@
Figure 2(c) shows that CCA, PPI++, and MI fail to achieve 95% nominal coverage (see Figure 2 (b)) primarily due to their biased estimation. For $\beta_{0}$ and $\beta_{1}$, WCCA results in negligible biases of 0.0005 and 0.0006, respectively. On the other hand, PS-PPI results in negligible biases of 0.0007 and 0.0008, even with the worst quality of ML predictions. For the parameters that CCA, PPI++, and MI fail to achieve 95% nominal coverage, they exhibit greater biases. Specifically, MI shows a bias of 0.115 for the estimation of $\beta_{2}$, while CCA results in a bias of -0.043 for $\beta_{1}$. In terms of PPI++, the bias for $\beta_{1}$ increases from -0.023 to -0.043 when the quality of ML predictions degrades.
Similar findings are observed from Figure S1, S2, S3 in the Supplementary Materials. The simulation results demonstrate that under non-monotone missingness patterns with the MAR assumption for the missing mechanism, the proposed PS-PPI approach can produce consistent estimates and valid inference with efficiency dominance over WCCA when the propensity score model is correctly estimated or known. Such consistency and validity in inference hold regardless of the quality of the ML predictions. Furthermore, when precise ML predictions are available, PS-PPI is much more efficient than WCCA. This corroborates our theoretical results.
@@
The simulation results under misspecified propensity score models are shown in Figure S4 and S5 in the Supplementary Materials. Note that WCCA and PS-PPI are the only two methods that are impacted by the propensity score model specification. When the propensity score model is misspecified in terms of the covariates, all the methods show a degree of performance in the width of the 95% CIs similar to the one shown in Figure 2. In terms of the coverage rate, Figure S4 and S5 show that WCCA fails to achieve the nominal 95% coverage rate for $\beta_{0}$ and $\beta_{1}$. PS-PPI achieves the 95% coverage rate when the ML predictions are sufficiently precise, while these coverage rates decline when the quality of ML predictions degrades. However, it can be seen that the coverage rates from PS-PPI still consistently outperform the WCCA baseline. Furthermore, PS-PPI also maintains a higher coverage rate and narrower intervals compared to the PPI++ method, which only utilizes the fully-observed and missing-outcome subsets of data.
5 Application to the All of Us Data
Background
Education attainment is a key social determinant of cardiovascular health, shaping lifestyle behaviors, access to healthcare, and long-term disease outcomes (wang2024association). Among the biological pathways linking social conditions to cardiovascular risk, systemic inflammation (often measured through C-reactive protein (CRP)) plays a central role. CRP has been widely associated with the development of cardiovascular disease (CVD)(ridker2003clinical). In this section, we use data from the All of Us (AoU) program (all2019all) to examine how education levels relate to CRP concentrations, accounting for demographic and clinical covariates, to explore whether disparities in inflammation help explain the socioeconomic gradient in CVD risk (magnani2024educational). AoU was launched in 2018 to establish a large-scale database of electronic health records (EHRs) and biospecimens for hypothesis-free population research. In this study, we analyze data from 314,625 participants who self-reported as male or female and have both EHR and genotype data available in the Curated Data Repository (CDR) version 8.
Analysis model with ML predictions
To investigate the association, we fit the following regression model with the primary covariate (education), adjusting for age, sex, race, smoking status, and the first 5 genetic principal components (PCs):
$\text{CRP}\sim$ $\text{Intercept}+\text{Education}+\text{Age}+\text{Sex}+\text{Race}+\text{Smoking}+\text{PC1}+\cdots+\text{PC5}.$ (11)
To address genetic relatedness, we keep one individual from each kinship group for analysis and use the rest to train ML models (mccaw2024synthetic); this results in the total $314,625$ split into i) 20,278 participants for ML model building, and ii) 294,347 participants for the analytic sample. In the analytical sample, 47,317 subjects are fully observed; Missingness occurs in race, smoking status, education, and CRP (15 patterns with at least one missing entry). Besides the complete cases, only patterns with more than 10,000 samples are kept: 1) missing CRP only (172,685); 2) missing race and CRP (50,040); 3) missing race only (10,934). We drop the remaining patterns with a total of 13,371 samples (4.54% of the analytical sample). This analysis decision simplifies modeling of the probabilities of missingness for 4 instead of 16 patterns with a significant portion of the original analytic sample retained for analysis. Each missing variable is imputed using either gradient boosting decision trees (XGBoost, chen2016xgboost) or random forest, whichever performs better on a validation set. Predictors for imputation include age, sex, genetically inferred ancestry, PCs, and diagnostic events coded by Systematized Medical Nomenclature for Medicine–Clinical Terminology (SNOMED) (stearns2001snomed). We apply the PS models in the form proposed by sun2018inverse (similar to Eq. (9) and (10)) and include all the available variables in each pattern for fitting. We compare against WCCA, or using only the missing outcome pattern for PS-PPI. We conduct sensitivity analysis of PS-PPI without IPW and compare against CCA and its outcome-based variant. For each method, we report coefficient estimates, standard error, 95% CIs, and variance ratio (relative to WCCA or CCA).
Table 1: Summary statistics for different subsets of the All of Us analytic sample used in this study: 1) complete cases ($C$); 2) $C$ + missing outcome only data; 3) $C$ + data with missingness patterns of size $\geqslant$ 10K; 4) all data. For C-reactive protein (CRP), mean, standard deviation, and range are reported. For other variables, the values represent percentages; counts are shown in parentheses. “Missing” rows show counts only. Missingness in smoking status and education levels are due to responses of “Skip”, “Don’t know”, or “Prefer not to answer”; missingness in race is due to responses of “Skip”, “None of these”, or “Prefer not to answer”. Note that age is treated as a continuous variable in the regression model.
$C$ CC+missing CRP CC+patterns of size $\geqslant$ 10K Total
Sample Size 47,317 220,002 280,976 294,347
C-reactive protein (CRP; mg/L)
Mean (SD) 3.9 (5.0) 3.9 (5.0) 4.1 (5.1) 4.1 (5.1)
Range 0.0–26.1 0.0–26.1 0.0–26.1 0.0–26.1
College graduate or higher 50.1 (23,729) 51.6 (113,617) 46.5 (130,688) 46.3 (133,523)
Missing (0) (0) (0) (6,072)
Age group
18–24 0.4 (170) 0.7 (1,464) 0.8 (2,312) 0.8 (2,489)
25–34 4.9 (2,332) 9.4 (20,575) 10.9 (30,497) 10.9 (31,998)
35–44 9.7 (4,583) 13.8 (30,386) 14.9 (41,829) 14.9 (43,882)
45–54 12.1 (5,734) 13.3 (29,193) 14.3 (40,168) 14.3 (42,150)
55–64 20.0 (9,468) 19.4 (42,609) 19.4 (54,384) 19.5 (57,277)
65+ 52.9 (25,030) 43.5 (95,775) 39.8 (111,786) 39.6 (116,551)
Sex at birth
Female 61.5 (29,102) 59.6 (131,070) 60.9 (171,010) 60.6 (178,463)
Male 38.5 (18,215) 40.4 (88,932) 39.1 (109,966) 39.4 (115,884)
Race
Asian 2.3 (1,083) 3.9 (8,666) 3.9 (8,666) 3.9 (8,995)
White 78.8 (37,273) 72.7 (159,848) 72.7 (159,848) 71.8 (164,883)
Black 16.6 (7,862) 20.9 (45,993) 20.9 (45,993) 21.7 (49,708)
Others 2.3 (1,099) 2.5 (5,495) 2.5 (5,495) 2.6 (5,968)
Missing (0) (0) (60,974) (64,793)
Have ever smoked $\geqslant$100 cigarettes 45.4 (21,463) 42.1 (92,627) 40.3 (113,103) 40.3 (115,628)
Missing (0) (0) (0) (7,731)
Depression 42.1 (19,932) 26.5 (58,217) 26.2 (73,746) 26.2 (77,196)
Hypertension 63.2 (29,909) 42.4 (93,384) 41.3 (115,996) 41.2 (121,409)
Diabetes 30.6 (14,467) 18.5 (40,739) 19.5 (54,732) 19.5 (57,469)
Anxiety 45.6 (21,596) 29.0 (63,872) 28.6 (80,450) 28.6 (84,152)
Obesity 45.0 (21,289) 28.8 (63,268) 29.5 (82,849) 29.4 (86,515)
Table 2: All of Us based coefficient estimates and 95% CIs for education level in the regression model in Eq. (11). Smaller variance ratios represent larger efficiency gains.
Method $\widehat{\beta}_{\textrm{edu}}$ Standard Error 95% CIs Variance Ratio
without IPW CCA (baseline) -1.187 0.048 [1.281,1.093][-1.281,-1.093] 1
PS-PPI (outcome missing only) -0.946 0.045 [1.035,0.857][-1.035,-0.857] 0.888
PS-PPI -1.091 0.042 [1.174,1.008][-1.174,-1.008] 0.778
with IPW WCCA (baseline) -1.192 0.053 [1.295,1.088][-1.295,-1.088] 1
PS-PPI (outcome missing only) -0.928 0.050 [1.027,0.830][-1.027,-0.830] 0.903
PS-PPI -1.110 0.047 [1.202,1.017][-1.202,-1.017] 0.796
Results
Table 1 shows the summary statistics of the included variables and several key conditions across different data subsets. Specifically, we compute the summary statistics for 1) the complete-case subset; 2) the subset consisting of complete cases and outcome-missing-only cases; 3) the subset of complete cases and all missingness patterns with more than 10,000 samples; and 4) the full dataset. The characteristics of several variables, including both the outcome and the covariate of interest, differ across these subsets. Specifically, in the subset with only complete cases and outcome-missing cases, the sample mean of the CRP level (in mg/L) is 3.9, compared with 4.1 in the full data. Similarly, the proportion of participants with a college degree or higher is 50.1% and 51.6% for the complete-case and complete-case plus outcome-missing subsets, respectively, both higher than that in the full dataset (46.3%). These discrepancies indicate that CCA and outcome-based prediction-based inference methods rely on subsets with distributions shifted from the overall population, thereby making the result less generalizable to the target population. Even if a method that uses the cases with complete or outcome-missing-only data can be modified to account for such distributional differences using propensity scores for outcome missingness, participants with other missingness patterns and the associated information are not used. In contrast, for the subset including complete cases and all patterns with more than 10,000 samples, the sample mean of CRP (4.1) matches that of the full data, and the proportion of college graduates (46.5%) closely approximates the full-sample value (46.3%). This demonstrates that the proposed PS-PPI method enables inference based on data subsets that are more representative of the full data distribution, thereby improving the generalizability of the inferential results under data missingness.
In terms of the statistical analysis, from Table 2, all methods show that a higher education level is significantly associated with a lower CRP level, adjusting for age, sex, race, smoking status, and the first 5 genetic PCs. Specifically, the WCCA estimate for ${\beta}_{\textrm{edu}}$ is 1.192-1.192 (95% CI: [-1.295, -1.088]). After integrating additional data from the three retained missingness patterns (each of size $\geqslant$ 10K) and ML imputations, the PS-PPI achieves a 20.4% variance reduction relative to WCCA, resulting in an estimate of 1.110-1.110 (95 % CI: [-1.202, -1.017]). We observe a smaller magnitude of variance reduction (9.7%) when integrating additional data but with only CRP missing and predicted. As a sensitivity analysis, without IPW, the CCA estimate is 1.187-1.187 (95% CI: [-1.281, -1.093]), similar to the WCCA estimate, indicating a small degree of selection bias of having CRP test results. The percent variance reduction of PS-PPI with only complete plus missing CRP data and PS-PPI with data from the three retained missingness patterns is 11.2% and 22.2%, respectively, favoring PS-PPI, which integrates more data and ML predictions.